• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name hsmith

    Log entry time 14:58:42 on May 02, 2011

    Entry number 357827

    keyword=Q2 Quench

    The left Q2 quenched around 8:19 am this morning. However, after carefully analyzing the data from the archive no exceptions as to what caused the quench could be found. Had to manually reset the left Q3, liquid level and lead flow were OK, no known issues found